OpenStudy (anonymous):

You divide two monomials and still have a fraction. Can this be correct? Explain.

OpenStudy (anonymous):

@whpalmer4

OpenStudy (whpalmer4):

Sure. If the numerator monomial isn't a multiple of the denominator monomial, you'll end up with a fraction.

OpenStudy (anonymous):

Thanks :)

OpenStudy (whpalmer4):

You can test this, of course: \[\frac{x+1}{x} = 1 + \frac{1}{x}\]If we had tried it with "real numbers", say \(x=2\), we'd get \[\frac{2+1}{2} = \frac{3}{2} = 1 + \frac{1}{2}\]and that comes as no surprise, right?

OpenStudy (whpalmer4):

If we make \(x\) bigger and bigger, the remainder fraction gets smaller and smaller, but it never quite becomes 0.